You are about to take a test that contains computation problems worth 4 points each and word problems worth 7 points each. You can do a computation problem in 2 minutes and a word problem in 4 minutes. You have 48 minutes to take the test and may answer no more than 14 problems. Assume that you answer all the problems attempted correctly.
Answer by ankor@dixie-net.com(22740)

:
You are about to take a test that contains computation problems worth 4 points each and word problems worth 7 points each.
You can do a computation problem in 2 minutes and a word problem in 4 minutes. You have 48 minutes to take the test and may answer no more than 14 problems.
Assume that you answer all the problems attempted correctly.
:
Let w = no. of word problems
Let c = no. of Computation problems
:
Time equation
2c + 4w = 48
:
No. of problems equation
c + w = 14
multiply above equation by -2, add to the 1st equation
2c + 4w = 48
-2c -2w = -28
---------------addition eliminates c, find w
2w = 20
w = 10 word problems
then
14-10 = 4 computation problems